OpenStudy (anonymous):

f(n+1)=1/2{f(n)+9/f(n)},n belongs to natural nos. and f(n)>0 for all n belonging to natural nos. then limit n tends to infinity equals to what

OpenStudy (phi):

if the limit exists, then f(n+1) approaches f(n). call this value x then x= (1/2) (x+9/x) 2x= x+9/x x=9/x x^2= 9 x= sqrt(9)

OpenStudy (anonymous):

thank you

OpenStudy (phi):

f(n)>0 for all n so it looks like x= +3 is the limit

OpenStudy (anonymous):

the answer is correct

OpenStudy (anonymous):

@phi-can you help me with a question from complex numbers

OpenStudy (phi):

?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f621417e4b079c5c6312ef5 this is the link to the question